# A Unified Linear Programming Framework for Offline Reward Learning from Human Demonstrations and Feedback

### v1

@@ Appendix D Comparison to Pessimism-based Approach
The concurrent work by [25] proposed an offline IRL algorithm for finite-horizon MDPs with comparable sample complexity, based on pessimistic value iteration. To be specific, they recover the mapping from the value and advantage functions to reward functions through Bellman iterations under estimated state transition probabilities. Though a direct comparison is limited since our work is developed for infinite-horizon discounted MDPs, there are some common structures between their algorithm and ours. Specifically, the value and advantage functions in their work can be considered as the primal optimization variable and the slack in the primal feasibility constraint in our formulation.
@@
Appendix E Proof of Theorem 2
@@
$g(\tau^{1},\tau^{2},y):=\mathbf{1}_{\{\mathcal{L}(\tau^{1},\tau^{2},y;r)\geq\epsilon_{r}\}}(\tau^{1},\tau^{2},y),$ (E.1)
@@
$\bar{g}=\mathbb{E}_{(\tau^{1},\tau^{2},y)\sim(\mu_{\text{HF}},\Phi)}[g(\tau^{1},\tau^{2},y)]=\mathbb{P}(\mathcal{L}(\tau^{1},\tau^{2},y;r)\geq\epsilon_{r}).$ (E.2)
@@ Appendix F Robustness of LP Framework
$\hat{\theta}_{\text{MLE}}\in\operatorname*{arg\,max}_{\theta\in\Theta}\sum_{n=1}^{N_{q}}\log\left(\Phi\left(-\mathcal{L}(\tau^{n,1},\tau^{n,2},y^{n};r_{\theta})\right)\right).$ (F.1)
@@
$\hat{\pi}_{\text{LP}}\in\operatorname*{arg\,max}_{\pi}\mathbb{E}_{s\sim d^{\pi}}[r_{\hat{\theta}_{\text{LP}}}(s,\pi(s))],\quad\hat{\pi}_{\text{MLE}}\in\operatorname*{arg\,max}_{\pi}\mathbb{E}_{s\sim d^{\pi}}[r_{\hat{\theta}_{\text{MLE}}}(s,\pi(s))].$ (F.2)
@@
Recent works in offline RLHF, such as those by [20, 19] have adapted the concept of pessimism from offline RL theory to address the data coverage issue. Specifically, these studies define a confidence set for the reward function and solve robust optimization problem to identify the policy that maximizes the worst-case reward within this set. For example, [19] uses a semi-norm $\|\cdot\|_{\Sigma+\lambda I}$ as a metric for constructing the confidence set, where $\Sigma$ represents the covariance of the comparison data and $\lambda>0$ is a conservativeness parameter, such that
$\mathcal{D}_{\text{PE}}=\left\{\theta\in\Theta\mid\|\theta-\hat{\theta}_{\text{MLE}}\|_{\Sigma+\lambda I}\leq f(N,k,\delta,\lambda)\right\}.$ (F.3)
@@
$\hat{\pi}_{\text{PE}}\in\operatorname*{arg\,max}_{\pi}\min_{\theta\in\mathcal{D}_{\text{PE}}}\mathbb{E}_{s\sim d^{\pi}}[r_{\theta}(s,\pi(s))].$ (F.4)
[19] prove that the true parameter $\theta^{*}$ exists in this confidence set with high probability as the number of sample increases, which enables them to provide an optimality guarantee.
@@
We construct a sampling distribution such that both MLE and pessimistic MLE algorithms in [19] returns a wrong policy with high probability, while LP succeeds to find an optimal policy. Specifically, if the pair $(a_{1},a_{2})$ is sampled with a significantly higher probability compared to the pair $(a_{2},a_{3})$ in the queries, we show that $\{\hat{\theta}_{\text{MLE}}\}_{2}>\{\hat{\theta}_{\text{MLE}}\}_{3}$ holds under the BTL model and the greedy evaluator. The MLE algorithm proposed in [19] estimate the reward parameter by solving
$\hat{\theta}_{\text{MLE}}\in\operatorname*{arg\,max}_{\theta\in\Theta}\frac{N_{12}}{N}\log\frac{e^{\theta_{2}}}{e^{\theta_{1}}+e^{\theta_{2}}}+\frac{N_{23}}{N}\log\frac{e^{\theta_{3}}}{e^{\theta_{2}}+e^{\theta_{3}}}+\frac{N_{31}}{N}\log\frac{e^{\theta_{3}}}{e^{\theta_{3}}+e^{\theta_{1}}},$ (F.5)
@@
$\begin{split}&\max_{\theta\in\Theta}\;J(\theta_{1},\theta_{2},\theta_{3})=\alpha\log\frac{e^{\theta_{2}}}{e^{\theta_{1}}+e^{\theta_{2}}}+\beta\log\frac{e^{\theta_{3}}}{e^{\theta_{2}}+e^{\theta_{3}}}\\ \end{split}$ (F.6)
@@
$\begin{split}\left[\frac{\partial\mathcal{L}}{\partial\theta_{2}}\right]_{(\theta^{*},\lambda^{*})}&=\alpha\frac{e^{\theta^{*}_{1}}}{e^{\theta^{*}_{1}}+e^{\theta^{*}_{2}}}-\beta\frac{e^{\theta^{*}_{2}}}{e^{\theta^{*}_{2}}+e^{\theta^{*}_{3}}}-2\lambda^{*}_{1}\theta^{*}_{2}+\lambda^{*}_{2}=0,\\ \left[\frac{\partial\mathcal{L}}{\partial\theta_{3}}\right]_{(\theta^{*},\lambda^{*})}&=\beta\frac{e^{\theta^{*}_{2}}}{e^{\theta^{*}_{2}}+e^{\theta^{*}_{3}}}-2\lambda^{*}_{1}\theta^{*}_{3}+\lambda^{*}_{2}=0.\end{split}$ (F.7)
@@
$\alpha\frac{e^{\theta^{*}_{1}}}{e^{\theta^{*}_{1}}+e^{\theta^{*}_{2}}}-2\beta\frac{e^{\theta^{*}_{2}}}{e^{\theta^{*}_{2}}+e^{\theta^{*}_{3}}}=2\lambda^{*}_{1}(\theta^{*}_{2}-\theta^{*}_{3}).$ (F.8)
@@
$\begin{split}\alpha\frac{e^{\theta^{*}_{1}}}{e^{\theta^{*}_{1}}+e^{\theta^{*}_{2}}}-2\beta\frac{e^{\theta^{*}_{2}}}{e^{\theta^{*}_{2}}+e^{\theta^{*}_{3}}}&>2\beta\left(\frac{e^{\theta^{*}_{1}+3}}{e^{\theta^{*}_{1}}+e^{\theta^{*}_{2}}}-\frac{e^{\theta^{*}_{2}}}{e^{\theta^{*}_{2}}+e^{\theta^{*}_{3}}}\right)=2\beta\frac{e^{\theta^{*}_{1}+\theta^{*}_{2}+3}+e^{\theta^{*}_{1}+\theta^{*}_{3}+3}-e^{\theta^{*}_{1}+\theta^{*}_{2}}-e^{2\theta^{*}_{2}}}{(e^{\theta^{*}_{1}}+e^{\theta^{*}_{2}})(e^{\theta^{*}_{2}}+e^{\theta^{*}_{3}})}\\ &=2\beta\frac{(e^{\theta^{*}_{1}+\theta^{*}_{2}+3}-e^{\theta^{*}_{1}+\theta^{*}_{2}})+(e^{3-\theta^{*}_{2}}-e^{2\theta^{*}_{2}})}{(e^{\theta^{*}_{1}}+e^{\theta^{*}_{2}})(e^{\theta^{*}_{2}}+e^{\theta^{*}_{3}})}>0,\end{split}$ (F.9)
@@
By Lemma 3, there exist $\alpha,\beta\in(0,1)$ such that if $\frac{N_{12}}{N}\geq\alpha$, $\frac{N_{23}}{N}\leq\beta$, and $N_{31}=0$, then $\{\hat{\theta}_{\text{MLE}}\}_{2}>\{\hat{\theta}_{\text{MLE}}\}_{3}$. For any $\delta>0$, there exists a sampling distribution $\mu_{\text{HF}}$ satisfying
@@
Next, we consider the pessimistic MLE under $\mu_{\text{HF}}$. The pessimistic MLE imposes higher penalty on the reward function of state-action pairs that have less support in the data. Therefore, intuitively, the penalty for the state $a_{3}$ will be higher than $a_{2}$, and thus, $\hat{\pi}_{\text{PE}}(s)=a_{2}$ will hold. We use the penalty function proposed in [19] to confirm this. Consider the covariance matrix
$\Sigma=\frac{1}{N}\begin{bmatrix}N_{12}&-N_{12}&0\\ -N_{12}&N_{12}+N_{23}&-N_{23}\\ 0&-N_{23}&N_{23}\end{bmatrix}.$ (F.11)
@@
$\begin{split}\phi_{2}&=\|[0,1,0]\|_{(\Sigma+\lambda I)^{-1}}^{2}=\frac{(N_{12}+\lambda)(N_{23}+\lambda)}{|\Sigma+\lambda I|},\\ \phi_{3}&=\|[0,0,1]\|_{(\Sigma+\lambda I)^{-1}}^{2}=\frac{(N_{12}+\lambda)(N_{12}+N_{23}+\lambda)-N_{12}^{2}}{|\Sigma+\lambda I|}.\end{split}$ (F.12)
@@
$\{\hat{\theta}_{\text{MLE}}\}_{2}-c\|\phi_{2}\|_{(\Sigma+\lambda I)^{-1}}>\{\hat{\theta}_{\text{MLE}}\}_{3}-c\|\phi_{3}\|_{(\Sigma+\lambda I)^{-1}}$ (F.13)
@@ Appendix G Extension to Continuous Feedback
$\mathbb{P}(y\leq\alpha\mid(\tau^{1},\tau^{2}))=\Phi(\alpha;r_{\text{true}}^{\top}(\psi^{1}-\psi^{2}))\quad\forall\alpha\in[-1,1].$ (G.1)
@@
$\mathcal{L}^{\prime}(\tau^{n,1},\tau^{n,2},y^{n};r):=\left(cy^{n}+r^{\top}(\psi^{n,2}-\psi^{n,1})\right)\mathbf{1}_{\{y^{n}\geq 0\}}+\left(-cy^{n}+r^{\top}(\psi^{n,1}-\psi^{n,2})\right)\mathbf{1}_{\{y^{n}\leq 0\}},$ (G.2)
where $c>0$ is a scaling parameter. The solution set ^$\hat{\mathcal{R}}_{\text{CHF}}(\epsilon_{r})$ is then defined in the same way with (4.3) using the error $\mathcal{L}^{\prime}$.
^$\hat{\mathcal{R}}_{\text{CHF}}(\epsilon_{r}):=\{r\mid\mathcal{L}^{\prime}(\tau^{n,1},\tau^{n,2},y^{n};r)\leq\epsilon_{r}\quad\forall n=1,2,\ldots,N_{q},\quad r\in[-1,1]^{|S||A|}\}.$ (G.3)
@@
^$\hat{\mathcal{R}}_{\text{IRL-CHF}}(\epsilon_{g},\epsilon_{r})=\{(r,u,v)\mid(u,v)\in\hat{\mathcal{R}}_{\text{IRL}}(\epsilon_{g}),\;r\in\hat{\mathcal{R}}_{\text{CHF}}(\epsilon_{r}),\;u=\hat{d}_{e}\circ r\}.$ (G.4)
It is important to note that ^$\hat{\mathcal{R}}_{\text{IRL-CHF}}(\epsilon_{g},\epsilon_{r})$ can become infeasible if $\epsilon_{g}$ or $\epsilon_{r}$ is set too small, due to the strict reward gap. Therefore, choosing proper values for $\epsilon_{g}$ and $\epsilon_{r}$ is crucial to ensure the feasibility of the LP. The generalization guarantee provided in Theorem 2 also holds with $\mathcal{L}^{\prime}$, under a similar proof.
@@ Appendix H Detailed Experimental Setup
Table 1: Algorithm Details
Algorithms Description Parameters
MLE-IRL Bi-level optimization algorithm for offline IRL [18] Step size $=0.01$
LP-IRL-1 LP formulation of IRL (3.20) with a moderate $\epsilon_{g}$ $\epsilon_{g}=0.01/\sqrt{N}$
LP-IRL-2 LP formulation of IRL (3.20) with a tighter $\epsilon_{g}$ $\epsilon_{g}=0.001/\sqrt{N}$
LP-IRL-D Integration of IRL and RLHF with discrete feedback (4.5) $\epsilon_{g}=0.01/\sqrt{N}$
$\epsilon_{r}=0.01/\sqrt{N}$
LP-IRL-C Integration of IRL and RLHF with continuous feedback (G.4) $\epsilon_{g}=0.1/\sqrt{N}$
$\epsilon_{r}=0.01/\sqrt{N}$
MLE-HF Pessimistic MLE under the BTL model [19] $\lambda=0.1$, $B=1$
LP-HF LP formulation of RLHF (4.3) $\epsilon_{r}=-0.01$
Table 1 provides a detailed description of the algorithms used in experiments. We employ a tabular setting for the reward function without any function approximation. In LP-IRL-1 and LP-IRL-2, we assume that $2/3$ of trajectories are sampled from $\pi_{e}$, and the remaining samples are obtained from the uniform policy to estimate $\hat{d}_{\text{sub}}$. In LP-IRL-D and LP-IRL-C, we assume that $2/3$ of trajectories are sampled from $\pi_{e}$, and the the remaining trajectories are sampled from the uniform policy to generate human feedback. In all LP-IRL algorithms, we use the $\mathcal{L}^{1}$ norm constraint for $X$, and we set $\delta=0.1$ and $B=100$. In MLE-HF and LP-HF, human feedback is given as discrete value following the greedy model. In LP-HF, a reward function is selected from (4.3) by optimizing a dummy objective function. For algorithm details of MLE-IRL and MLE-RLHF, please refer to [18] and [19], respectively.

### v3

@@ Appendix D Comparison to Pessimism-based Approach
The concurrent work by (Zhao et al., 2023) proposed an offline IRL algorithm for finite-horizon MDPs with comparable sample complexity, based on pessimistic value iteration. To be specific, they recover the mapping from the value and advantage functions to reward functions through Bellman iterations under estimated state transition probabilities. Though a direct comparison is limited since our work is developed for infinite-horizon discounted MDPs, there are some common structures between their algorithm and ours. Specifically, the value and advantage functions in their work can be considered as the primal optimization variable and the slack in the primal feasibility constraint in our formulation.
@@
Appendix E Proof of Theorem 4.1
@@
$g(\tau^{1},\tau^{2},y):=\mathbf{1}_{\{\mathcal{L}(\tau^{1},\tau^{2},y;r)\geq\epsilon_{r}\}}(\tau^{1},\tau^{2},y),$ (56)
@@
$\bar{g}=\mathbb{E}_{(\tau^{1},\tau^{2},y)\sim(\mu_{\text{HF}},\Phi)}[g(\tau^{1},\tau^{2},y)]=\mathbb{P}(\mathcal{L}(\tau^{1},\tau^{2},y;r)\geq\epsilon_{r}).$ (57)
@@ Appendix F Robustness of LP Framework
$\hat{\theta}_{\text{MLE}}\in\operatorname*{arg\,max}_{\theta\in\Theta}\sum_{n=1}^{N_{q}}\log\left(\Phi\left(-\mathcal{L}(\tau^{n,1},\tau^{n,2},y^{n};r_{\theta})\right)\right).$ (59)
@@
$\hat{\pi}_{\text{LP}}\in\operatorname*{arg\,max}_{\pi}\mathbb{E}_{s\sim d^{\pi}}[r_{\hat{\theta}_{\text{LP}}}(s,\pi(s))],\quad\hat{\pi}_{\text{MLE}}\in\operatorname*{arg\,max}_{\pi}\mathbb{E}_{s\sim d^{\pi}}[r_{\hat{\theta}_{\text{MLE}}}(s,\pi(s))].$ (60)
@@
Recent works in offline RLHF, such as those by (Zhan et al., 2023; Zhu et al., 2023) have adapted the concept of pessimism from offline RL theory to address the data coverage issue. Specifically, these studies define a confidence set for the reward function and solve robust optimization problem to identify the policy that maximizes the worst-case reward within this set. For example, (Zhu et al., 2023) uses a semi-norm $\|\cdot\|_{\Sigma+\lambda I}$ as a metric for constructing the confidence set, where $\Sigma$ represents the covariance of the comparison data and $\lambda>0$ is a conservativeness parameter, such that
$\mathcal{D}_{\text{PE}}=\left\{\theta\in\Theta\mid\|\theta-\hat{\theta}_{\text{MLE}}\|_{\Sigma+\lambda I}\leq f(N,k,\delta,\lambda)\right\}.$ (61)
@@
$\hat{\pi}_{\text{PE}}\in\operatorname*{arg\,max}_{\pi}\min_{\theta\in\mathcal{D}_{\text{PE}}}\mathbb{E}_{s\sim d^{\pi}}[r_{\theta}(s,\pi(s))].$ (62)
(Zhu et al., 2023) prove that the true parameter $\theta^{*}$ exists in this confidence set with high probability as the number of sample increases, which enables them to provide an optimality guarantee.
@@
We construct a sampling distribution such that both MLE and pessimistic MLE algorithms in (Zhu et al., 2023) returns a wrong policy with high probability, while LP succeeds to find an optimal policy. Specifically, if the pair $(a_{1},a_{2})$ is sampled with a significantly higher probability compared to the pair $(a_{2},a_{3})$ in the queries, we show that $\{\hat{\theta}_{\text{MLE}}\}_{2}>\{\hat{\theta}_{\text{MLE}}\}_{3}$ holds under the BTL model and the greedy evaluator. The MLE algorithm proposed in (Zhu et al., 2023) estimate the reward parameter by solving
$\hat{\theta}_{\text{MLE}}\in\operatorname*{arg\,max}_{\theta\in\Theta}\frac{N_{12}}{N}\log\frac{e^{\theta_{2}}}{e^{\theta_{1}}+e^{\theta_{2}}}+\frac{N_{23}}{N}\log\frac{e^{\theta_{3}}}{e^{\theta_{2}}+e^{\theta_{3}}}+\frac{N_{31}}{N}\log\frac{e^{\theta_{3}}}{e^{\theta_{3}}+e^{\theta_{1}}},$ (63)
@@
$\begin{split}&\max_{\theta\in\Theta}\;J(\theta_{1},\theta_{2},\theta_{3})=\alpha\log\frac{e^{\theta_{2}}}{e^{\theta_{1}}+e^{\theta_{2}}}+\beta\log\frac{e^{\theta_{3}}}{e^{\theta_{2}}+e^{\theta_{3}}}\\ \end{split}$ (64)
@@
$\begin{split}\left[\frac{\partial\mathcal{L}}{\partial\theta_{2}}\right]_{(\theta^{*},\lambda^{*})}&=\alpha\frac{e^{\theta^{*}_{1}}}{e^{\theta^{*}_{1}}+e^{\theta^{*}_{2}}}-\beta\frac{e^{\theta^{*}_{2}}}{e^{\theta^{*}_{2}}+e^{\theta^{*}_{3}}}-2\lambda^{*}_{1}\theta^{*}_{2}+\lambda^{*}_{2}=0,\\ \left[\frac{\partial\mathcal{L}}{\partial\theta_{3}}\right]_{(\theta^{*},\lambda^{*})}&=\beta\frac{e^{\theta^{*}_{2}}}{e^{\theta^{*}_{2}}+e^{\theta^{*}_{3}}}-2\lambda^{*}_{1}\theta^{*}_{3}+\lambda^{*}_{2}=0.\end{split}$ (65)
@@
$\alpha\frac{e^{\theta^{*}_{1}}}{e^{\theta^{*}_{1}}+e^{\theta^{*}_{2}}}-2\beta\frac{e^{\theta^{*}_{2}}}{e^{\theta^{*}_{2}}+e^{\theta^{*}_{3}}}=2\lambda^{*}_{1}(\theta^{*}_{2}-\theta^{*}_{3}).$ (66)
@@
$\begin{split}\alpha\frac{e^{\theta^{*}_{1}}}{e^{\theta^{*}_{1}}+e^{\theta^{*}_{2}}}-2\beta\frac{e^{\theta^{*}_{2}}}{e^{\theta^{*}_{2}}+e^{\theta^{*}_{3}}}&>2\beta\left(\frac{e^{\theta^{*}_{1}+3}}{e^{\theta^{*}_{1}}+e^{\theta^{*}_{2}}}-\frac{e^{\theta^{*}_{2}}}{e^{\theta^{*}_{2}}+e^{\theta^{*}_{3}}}\right)=2\beta\frac{e^{\theta^{*}_{1}+\theta^{*}_{2}+3}+e^{\theta^{*}_{1}+\theta^{*}_{3}+3}-e^{\theta^{*}_{1}+\theta^{*}_{2}}-e^{2\theta^{*}_{2}}}{(e^{\theta^{*}_{1}}+e^{\theta^{*}_{2}})(e^{\theta^{*}_{2}}+e^{\theta^{*}_{3}})}\\ &=2\beta\frac{(e^{\theta^{*}_{1}+\theta^{*}_{2}+3}-e^{\theta^{*}_{1}+\theta^{*}_{2}})+(e^{3-\theta^{*}_{2}}-e^{2\theta^{*}_{2}})}{(e^{\theta^{*}_{1}}+e^{\theta^{*}_{2}})(e^{\theta^{*}_{2}}+e^{\theta^{*}_{3}})}>0,\end{split}$ (67)
@@
By Lemma F.2, there exist $\alpha,\beta\in(0,1)$ such that if $\frac{N_{12}}{N}\geq\alpha$, $\frac{N_{23}}{N}\leq\beta$, and $N_{31}=0$, then $\{\hat{\theta}_{\text{MLE}}\}_{2}>\{\hat{\theta}_{\text{MLE}}\}_{3}$. For any $\delta>0$, there exists a sampling distribution $\mu_{\text{HF}}$ satisfying
@@
Next, we consider the pessimistic MLE under $\mu_{\text{HF}}$. The pessimistic MLE imposes higher penalty on the reward function of state-action pairs that have less support in the data. Therefore, intuitively, the penalty for the state $a_{3}$ will be higher than $a_{2}$, and thus, $\hat{\pi}_{\text{PE}}(s)=a_{2}$ will hold. We use the penalty function proposed in (Zhu et al., 2023) to confirm this. Consider the covariance matrix
$\Sigma=\frac{1}{N}\begin{bmatrix}N_{12}&-N_{12}&0\\ -N_{12}&N_{12}+N_{23}&-N_{23}\\ 0&-N_{23}&N_{23}\end{bmatrix}.$ (69)
@@
$\phi_{2}=\|[0,1,0]\|_{(\Sigma+\lambda I)^{-1}}^{2}=\frac{(N_{12}+\lambda)(N_{23}+\lambda)}{|\Sigma+\lambda I|},\quad\phi_{3}=\|[0,0,1]\|_{(\Sigma+\lambda I)^{-1}}^{2}=\frac{(N_{12}+\lambda)(N_{12}+N_{23}+\lambda)-N_{12}^{2}}{|\Sigma+\lambda I|}.$ (70)
@@
$\{\hat{\theta}_{\text{MLE}}\}_{2}-c\|\phi_{2}\|_{(\Sigma+\lambda I)^{-1}}>\{\hat{\theta}_{\text{MLE}}\}_{3}-c\|\phi_{3}\|_{(\Sigma+\lambda I)^{-1}}$ (71)
@@ Appendix G Extension to Continuous Feedback
$\mathbb{P}(y\leq\alpha\mid(\tau^{1},\tau^{2}))=\Phi(\alpha;r_{\text{true}}^{\top}(\psi^{1}-\psi^{2}))\quad\forall\alpha\in[-1,1].$ (72)
@@
$\mathcal{L}^{\prime}(\tau^{n,1},\tau^{n,2},y^{n};r):=\left(cy^{n}+r^{\top}(\psi^{n,2}-\psi^{n,1})\right)\mathbf{1}_{\{y^{n}\geq 0\}}+\left(-cy^{n}+r^{\top}(\psi^{n,1}-\psi^{n,2})\right)\mathbf{1}_{\{y^{n}\leq 0\}},$ (73)
where $c>0$ is a scaling parameter. The solution set ^$\hat{\mathcal{R}}_{\text{CHF}}(\epsilon_{r})$ is then defined in the same way with (29) using the error $\mathcal{L}^{\prime}$.
^$\hat{\mathcal{R}}_{\text{CHF}}(\epsilon_{r}):=\{r\mid\mathcal{L}^{\prime}(\tau^{n,1},\tau^{n,2},y^{n};r)\leq\epsilon_{r}\quad\forall n=1,2,\ldots,N_{q},\quad r\in[-1,1]^{|S||A|}\}.$ (74)
@@
^$\hat{\mathcal{R}}_{\text{IRL-CHF}}(\epsilon_{g},\epsilon_{r})=\{(r,u,v)\mid(u,v)\in\hat{\mathcal{R}}_{\text{IRL}}(\epsilon_{g}),\;r\in\hat{\mathcal{R}}_{\text{CHF}}(\epsilon_{r}),\;u=\hat{d}_{e}\circ r\}.$ (75)
It is important to note that ^$\hat{\mathcal{R}}_{\text{IRL-CHF}}(\epsilon_{g},\epsilon_{r})$ can become infeasible if $\epsilon_{g}$ or $\epsilon_{r}$ is set too small, due to the strict reward gap. Therefore, choosing proper values for $\epsilon_{g}$ and $\epsilon_{r}$ is crucial to ensure the feasibility of the LP. The generalization guarantee provided in Theorem 4.1 also holds with $\mathcal{L}^{\prime}$, under a similar proof.
@@ Appendix H Detailed Experimental Setup
Table 1: Algorithm Details
Algorithms Description Parameters
MLE-IRL Bi-level optimization algorithm for offline IRL (Zeng et al., 2023) Step size $=0.01$
LP-IRL-1 LP formulation of IRL (26) with a moderate $\epsilon_{g}$ $\epsilon_{g}=0.01/\sqrt{N}$
LP-IRL-2 LP formulation of IRL (26) with a tighter $\epsilon_{g}$ $\epsilon_{g}=0.001/\sqrt{N}$
LP-IRL-D Integration of IRL and RLHF with discrete feedback (31) $\epsilon_{g}=0.01/\sqrt{N}$, $\epsilon_{r}=0.01/\sqrt{N}$
LP-IRL-C Integration of IRL and RLHF with continuous feedback (75) $\epsilon_{g}=0.1/\sqrt{N}$, $\epsilon_{r}=0.01/\sqrt{N}$
MLE-HF Pessimistic MLE under the BTL model (Zhu et al., 2023) $\lambda=0.1$, $B=1$
LP-HF LP formulation of RLHF (29) $\epsilon_{r}=-0.01$
@@
Table 1 provides a detailed description of the algorithms used in experiments. We employ a tabular setting for the reward function without any function approximation. In LP-IRL-1 and LP-IRL-2, we assume that $2/3$ of trajectories are sampled from $\pi_{e}$, and the remaining samples are obtained from the uniform policy to estimate $\hat{d}_{\text{sub}}$. In LP-IRL-D and LP-IRL-C, we assume that $2/3$ of trajectories are sampled from $\pi_{e}$, and the the remaining trajectories are sampled from the uniform policy to generate human feedback. In all LP-IRL algorithms, we use the $\mathcal{L}^{1}$ norm constraint for $X$, and we set $\delta=0.1$ and $B=100$. In MLE-HF and LP-HF, human feedback is given as discrete value following the greedy model. In LP-HF, a reward function is selected from (29) by optimizing a dummy objective function. For algorithm details of MLE-IRL and MLE-RLHF, please refer to (Zeng et al., 2023) and (Zhu et al., 2023), respectively.